Unicorn and Zebra form the UZ Co., a partnership. Unicorn contributes inventory with fair market value of $650,000 and an adjusted basis of $200,000 and cash of 350,000 for a 50% interest in the partnership. Zebra contributes computers, hardware, and proprietary software with a fair market value of $1,000,000 and a tax basis of $300,000 for a 50% interest in the partnership. The computer hardware, which was purchased from IBM on credit, is subject to an equipment loan of $300,000.

Please discuss the tax impact to:

  1. Unicorn,
  2. Zebra, and
  3. UZ Co.
Answer #1

Tax impact to Unicorn - The taxable income in the hands of Unicorn is to the extent of cash received over and above the adjusted tax basis = 350,000 - 200,000 = 150,000.

Tax impact to Zebra - Adjusted tax basis = 300,000 Equipment loan = 300,000 and interest in partnership = 50% Therefore, 50% of the loan is 150,000. Balance loan deductible from basis = 300,000 - 150,000 = 150,000

Therefore adjusted basis after deduction of above amount would be = 300,000 - 150,000 = 150,000

There will be no tax impact to Zebra in this case but the basis for computer hardware in the hands of the partnership will also be 150,000

Tax impact to UZ Co. - Adjusted basis of transfer from Unicorn and Zebra - Nil from Unicorn (as it is taxed in his hands) + 150,000 (as explained above) = 150,000